Determine the points of discontinuity of the function f and the nature of discontinuity at each of those points:


f ={-x², when x ≤ 0.


4-5x , when 0<x≤1


3x-4x², when 1<x≤2


-12x + 2x , when x>2}


Also check whether the function f is derivable at x = 1



1
Expert's answer
2022-03-15T19:22:29-0400

"f(x)= \begin{cases}
 -x² &\text{if } x≤0 \\
 4-5x&\text{if } 0<x≤1\\
 3x-4x²&\text{if } 1<x≤2\\
 -12x+2x &\text{if } x≥2
\end{cases}"


f(x) is polynomial function on every intervals (-∞,0), (0,1), (1,2) and (2,∞).

As polynomial function is always continuous function, here f(x) can have only discontinuites at x=0, 1, 2.

Let us check the continuity at these points.

At x=0

f(0-0) = "\lim_{x\to 0^-}{(-x²)}=0"

f(0+0) ="\lim_{x\to 0^+}{(4-5x)}=4"

f(0) = -0² = 0

Since f(0-0) ≠ f(0+0) , f(x) is discontinuous at x= 0

At x = 1

f(1-0) = "\lim_{x\to 1^-}{(4-5x)}=-1"

f(1+0) = "\lim_{x\to1^+}(3x-4x²)=-1"

f(1) = 4-5 = -1

Since f(1-0) =f(1)= f(1+0) , f(x) is continuous at x= 1


At x = 2

f(2-0) = "\lim_{x\to 2^-}{(3x-4x²)}=-10"

f(2+0) = "\lim_{x\to2^+}(-12x+2x)=-20"

f(2) = -24+4= -20

Since f(2-0)≠f(2+0), f(x) is discontinuous at x= 2

So points of discontinuity are x=0 and x=2